A sample of nitrogen gas expands in volume from 1.6 L to 5.4 L at constant temperature. Calculate the work done in joules if the gas expands against a constant pressure of 0.80 atm


1
Expert's answer
2021-09-18T23:58:32-0400

Work done , W = PdV

= P ×( Vf - Vi )

Where

P = external pressure = 0.80 atm = 0.80× 101.3×10^3 Pa( Since 1 atm = 101.3×10^3 Pa)

= 81040 Pa

Vf = final volume = 5.4 L

Vi = initial volume = 1.6 L

So dV = 5.4-1.6 = 3.8 L = 3.8×10^-3 m^3

So Work done ,W = 81040× 3.8×10^-3 J

= 307.9 J